# On thermal transpiration and thermomolecular pressure difference

### v1

@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.29)
@@
$\quad+C\frac{1}{\kappa}\Big(\displaystyle\int_{\{v_{1}\cdot n(q)>0\}\cap A}\int_{\frac{\mathcal{D}}{N}}^{\frac{\mathcal{D}}{|v_{1}|}}\int_{B}|k|(v_{1},u)f^{2}(q-tv_{1},u)w^{-2}(u)dudtdv_{1}\Big)^{1/2}$ (4.30)
@@
$\quad+C\frac{1}{\kappa}\Big(\displaystyle\int_{\Omega}\int_{\frac{\mathcal{D}}{N}}^{\infty}\int_{\mathbb{R}^{3}}\frac{1}{t^{3}}\nu(u)f^{2}(y,u)w^{-2}(u)dudtdy\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\left(\frac{1}{\kappa}\right)\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (4.31)
@@
$\lVert h^{l+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.33)
@@
Now, we construct the solution of (4.1) by starting from (4.2). It is clear that Property 4.1 still holds when $\hat{\alpha}=0$ or $\eta=1$. We rewrite 4.33 as
$\lVert h^{i+2(n-1)+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.34)
$\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.35)
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{i+m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\zeta}\left(\min\{1,\frac{1}{\kappa}\}\right)\max_{0\leq m\leq 2(n-1)}\lVert h^{i+1+m}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})},$
@@
$\lVert h^{i}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$ $=\lVert f^{i}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$ (4.36)
$\leq C_{\eta}\max\{\sqrt{\kappa},\kappa\}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
@@
$\max_{0\leq m\leq 2(n-1)}\lVert h^{i+m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\left(\frac{1}{8}\right)^{[\frac{i}{2n-1}]}\max_{0\leq m\leq 4(n-1)}\lVert h^{m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.37)
$\quad+C_{n,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
$\quad+C_{\eta}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big).$
@@
    $A_{i}^{k}\leq\left(\frac{1}{8}\right)^{[\frac{i}{k+1}]}\max\{A_{0}^{k},A_{1}^{k},\cdots,A_{k}^{k}\}+\frac{8+k}{7}D,\quad\text{for}\,i\leq k+1$ (4.38)
@@
    $A_{i}^{k}\leq\left(\frac{1}{8}\right)^{[\frac{i}{k+1}]}\max\{A_{0}^{k},A_{1}^{k},\cdots,A_{k}^{k}\}+2C_{k}\frac{8+k}{7}D,\quad\text{for}\,i\leq k+1$ (4.39)
@@
$\lVert h^{m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{n,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.40)
$\quad+C_{n,\eta}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
for $0\leq m\leq 4(n-1)$. Here, $C_{n,\eta},C_{n,\zeta}$ are increasing with respect to $n$. In particular, combining 4.37 and 4.40, we have
$\lVert h^{l}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\left(\frac{1}{8}\right)^{[\frac{l}{2n-1}]}C_{n,\eta,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.41)
$\qquad+C_{n,\eta,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big),$ (4.42)
@@
$\lVert h^{m_{1}+1}-h^{m_{2}+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{m_{1}-m}-h^{m_{2}-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.43)
$\qquad\qquad+C_{n}\min\{1,\frac{1}{\kappa}\}\max_{0\leq m\leq 2(n-1)}\lVert f^{m_{1}+1-m}-f^{m_{2}+1-m}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
@@
$C_{n}\max_{0\leq m\leq 2(n-1)}\lVert f^{m_{1}+1-m}-f^{m_{2}+1-m}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}<\frac{\epsilon}{2(2n+7)/7}.$ (4.44)
@@
$\lVert h^{m_{1}+1}-h^{m_{2}+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{m_{1}-m}-h^{m_{2}-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\epsilon.$ (4.45)
By (4.41), we find that
$\lVert h^{m_{1}}-h^{m_{2}}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{f_{in},\phi_{1},\phi_{2}}C_{n,\eta,\zeta}\left(\frac{1}{8}\right)^{[\frac{\min\{m_{1},m_{2}\}}{2n-1}]}+\epsilon/2.$ (4.46)
@@
$\lVert h^{\hat{\alpha},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (4.47)
$\qquad+C_{n}\left(\frac{1}{\kappa}\right)\lVert h^{\hat{\alpha},\eta}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
@@
$\qquad+C_{n,\eta}\left(\frac{1}{\kappa}\right)\lVert f^{\hat{\alpha},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{n,\eta,\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\max\{1,\kappa\}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
@@
$\lVert h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (4.48)
$\leq C_{\zeta}\kappa\max\{\hat{\alpha}_{1},\hat{\alpha}_{2}\}\left(\lVert\langle v\rangle^{-1}h^{\hat{\alpha}_{1},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\langle v\rangle^{-1}h^{\hat{\alpha}_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{n,\eta}\min\{1,\frac{1}{\kappa}\}\lVert f^{\hat{\alpha}_{1},\eta}-f^{\hat{\alpha}_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{\zeta}\kappa\max\{\hat{\alpha}_{1},\hat{\alpha}_{2}\}\left(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{n,\eta}\min\{1,\frac{1}{\kappa}\}\lVert f^{\hat{\alpha}_{1},\eta}-f^{\hat{\alpha}_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
@@
$\lVert h^{\eta_{1}}-h^{\eta_{2}}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{\zeta}|\eta_{1}-\eta_{2}|\cdot\lVert P_{w}(h^{\eta_{2}})\rVert_{L^{\infty}(\partial\Omega^{-}_{2})}+C_{n}\left(\frac{1}{\kappa}\right)\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$ (4.49)
$\leq C_{\zeta}|\eta_{1}-\eta_{2}|\cdot\lVert h^{\eta_{2}}\rVert_{L^{\infty}(\partial\Omega^{+})}+C_{n}\left(\frac{1}{\kappa}\right)\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
@@
$\qquad+C_{n}\min\{1,\frac{1}{\kappa}\}\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
The last inequality follows from a similar argument showing that $\lVert h^{\hat{\alpha},\eta}\rVert_{L^{\infty}}$ is bounded with respect to $\hat{\alpha},\eta$. Thus $\{h^{\eta}\}_{\eta}$ is a Cauchy sequence in $L^{\infty}$ and $h\coloneq\displaystyle\lim_{\eta\to 1}h^{\eta}$ is a mild solution of (4.1). The proof of the estimate (4.7) is similar to the proof of (4.47). ∎
@@
Now, we conclude this chapter by proving Theorem 1.1.
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}f_{k+1}+\frac{1}{\kappa}L(f_{k+1})=&\frac{1}{\kappa}\Gamma(f_{k},f_{k}),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f_{k+1}=&f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f_{k+1}=&P(f_{k+1})+m^{-1}(m_{T(x)}-m)P(f_{k})\\ &\quad+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (4.50)

### v2

@@ 4.1. The existence of solution to (80)
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (108)
@@
$\quad+\frac{C}{\kappa}\Big(\displaystyle\int_{\{v_{1}\cdot n(q)>0\}\cap A}\int_{\frac{\mathcal{D}}{N}}^{\frac{\mathcal{D}}{|v_{1}|}}\int_{B}e^{-\frac{C_{N}}{\kappa}t}|k|(v_{1},u)f^{2}(q-tv_{1},u)w^{-2}(u)dudtdv_{1}\Big)^{1/2}$ (109)
@@
$\quad+\frac{C}{\kappa}\Big(\displaystyle\int_{\Omega}\int_{\frac{\mathcal{D}}{N}}^{\infty}\int_{\mathbb{R}^{3}}\frac{1}{t^{3}}\nu(u)f^{2}(y,u)w^{-2}(u)dudtdy\Big)^{1/2}$
$\leq C\Big(\epsilon^{3-|\gamma|}_{1}+\frac{1}{1+N}\Big)\lVert f\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C\min\{1,\frac{1}{\kappa}\}\lVert f/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$ (110)
@@
$\lVert h^{l+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (112)
@@
Now, we construct the solution of (80) by starting from (81). It is clear that Property 4.1 still holds when $\hat{\alpha}=0$ or $\eta=1$. We rewrite (112) as
$\lVert h^{i+2(n-1)+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (113)
$\quad+\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{i+m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+C_{\zeta}\min\{1,\frac{1}{\kappa}\}\max_{0\leq m\leq 2(n-1)}\lVert h^{i+1+m}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})},$
@@
$\lVert h^{i}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$ $=\lVert f^{i}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$ (114)
$\leq C_{\eta}\max\{\kappa^{-1},\kappa\}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
@@
$\max_{0\leq m\leq 2(n-1)}\lVert h^{i+m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\left(\frac{1}{8}\right)^{[\frac{i}{2n-1}]}\max_{0\leq m\leq 4(n-1)}\lVert h^{m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (115)
$\quad+C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
$\quad+C_{\hat{\alpha},\eta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big).$
@@
    $A_{i}^{k}\leq\left(\frac{1}{8}\right)^{[\frac{i}{k+1}]}\max\{A_{0}^{k},A_{1}^{k},\cdots,A_{k}^{k}\}+\frac{8+k}{7}D,\quad\text{for}\,i\leq k+1$ (116)
@@
    $A_{i}^{k}\leq\left(\frac{1}{8}\right)^{[\frac{i}{k+1}]}\max\{A_{0}^{k},A_{1}^{k},\cdots,A_{k}^{k}\}+2C_{k}\frac{8+k}{7}D,\quad\text{for}\,i\leq k+1$ (117)
@@
$\lVert h^{m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (118)
$\quad+C_{\zeta,\hat{\alpha},\eta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{1}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}\Big),$
for $0\leq m\leq 4(n-1)$. Here, $C_{n,\eta},C_{n,\zeta}$ are increasing with respect to $n$. In particular, combining (115) and (118), we have
$\lVert h^{l}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\left(\frac{1}{8}\right)^{[\frac{l}{2n-1}]}C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (119)
$\qquad+C_{\zeta,\hat{\alpha},\eta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big),$
@@
$\lVert h^{m_{1}+1}-h^{m_{2}+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{m_{1}-m}-h^{m_{2}-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (120)
$\qquad\qquad+C_{\zeta}\min\{1,\kappa^{-1}\}\max_{0\leq m\leq 2(n-1)}\lVert f^{m_{1}+1-m}-f^{m_{2}+1-m}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
@@
$C_{\zeta}\max_{0\leq m\leq 2(n-1)}\lVert f^{m_{1}+1-m}-f^{m_{2}+1-m}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}<\frac{\epsilon}{2(2n+7)/7}.$ (121)
@@
$\lVert h^{m_{1}+1}-h^{m_{2}+1}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq\frac{1}{8}\max_{0\leq m\leq 2(n-1)}\lVert h^{m_{1}-m}-h^{m_{2}-m}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\epsilon.$ (122)
By (119), we find that
$\lVert h^{m_{1}}-h^{m_{2}}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{f_{in},\phi_{1},\phi_{2}}C_{n,\eta,\zeta}\left(\frac{1}{8}\right)^{[\frac{\min\{m_{1},m_{2}\}}{2n-1}]}+\epsilon/2.$ (123)
@@
$\lVert h^{\hat{\alpha},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$ (124)
$\qquad+\frac{C_{\zeta}}{\kappa}\lVert h^{\hat{\alpha},\eta}/w\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
@@
$\qquad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\hat{\alpha},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{\zeta}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\max\{1,\kappa\}\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
$\qquad+C_{\zeta,\hat{\alpha},\eta}\max\{1,\kappa^{-1}\}\Big(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\Big)$
@@
$\lVert h^{\hat{\alpha}_{1},\eta}-h^{\hat{\alpha}_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ (125)
$\leq C_{\zeta}\cdot\kappa\cdot\max\{\hat{\alpha}_{1},\hat{\alpha}_{2}\}\left(\lVert\langle v\rangle^{-1}h^{\hat{\alpha}_{1},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}+\langle v\rangle^{-1}h^{\hat{\alpha}_{2},\eta}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\hat{\alpha}_{1},\eta}-f^{\hat{\alpha}_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
$\leq C_{\zeta}\cdot\kappa\cdot\max\{\hat{\alpha}_{1},\hat{\alpha}_{2}\}\left(\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}+\kappa\lVert\phi_{1}\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\right)$
$\quad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\hat{\alpha}_{1},\eta}-f^{\hat{\alpha}_{2},\eta}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
@@
$\lVert h^{\eta_{1}}-h^{\eta_{2}}\rVert_{L^{\infty}(\Omega\times\mathbb{R}^{3})}$ $\leq C_{\zeta}|\eta_{1}-\eta_{2}|\cdot\lVert P_{w}(h^{\eta_{2}})\rVert_{L^{\infty}(\partial\Omega^{-}_{2})}+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$ (126)
$\leq C_{\zeta}|\eta_{1}-\eta_{2}|\cdot\lVert h^{\eta_{2}}\rVert_{L^{\infty}(\partial\Omega^{+})}+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}$
@@
$\qquad+C_{\zeta}\min\{1,\kappa^{-1}\}\lVert f^{\eta_{1}}-f^{\eta_{2}}\rVert_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}.$
The last inequality follows from a similar argument showing that $\lVert h^{\hat{\alpha},\eta}\rVert_{L^{\infty}}$ is bounded with respect to $\hat{\alpha},\eta$. Thus $\{h^{\eta}\}_{\eta}$ is a Cauchy sequence in $L^{\infty}$ and $h\coloneq\lim_{\eta\to 1}h^{\eta}$ is a mild solution of (80). The proof of the estimate (86) is similar to the proof of (124). ∎
@@
Now, we conclude this section by proving Theorem 1.1.
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}f_{k+1}+\frac{1}{\kappa}L(f_{k+1})=&\frac{1}{\kappa}\Gamma(f_{k},f_{k}),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f_{k+1}=&f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f_{k+1}=&P(f_{k+1})+m^{-1}(m_{T(x)}-m)P(f_{k})\\ &\quad+\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m),&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (127)